OpenStudy (anonymous):

using the side lengths, determone whether the triangle is acute, obtuse, or right

OpenStudy (anonymous):



OpenStudy (anonymous):

Let a=4,b=5,c=8 a^2+b^2<c^2 So, triangle is obtuse.

OpenStudy (anonymous):

16+64+25=105

OpenStudy (phi):

if your triangle was a right triangle with legs 4 and 5 its hypotenuse would be c^2 = 4^2 + 5^2 = 16+25= 41 c= sqrt(41) , or about 6.4